# A look at the operator product expansion in critical dynamics

### v1

@@ IV.1 Gaussian Fokker-Planck Hamiltonian
$\hat{\phi}\left(x\right)$ $=$ $\phi\left(x\right)$
$\hat{\pi}\left(x\right)$ $=$ $\frac{1}{i}\frac{\delta}{\delta\phi\left(x\right)}\,.$
@@
$\partial_{t}P\left[\phi\right]\left(t\right)$ $=$ $D\Biggr{\{}\int_{x}i\hat{\pi}\left(x\right)\left(-\partial_{x}^{2}+m^{2}\right)\hat{\phi}\left(x\right)-\int_{x}\hat{\pi}\left(x\right)^{2}\Biggr{\}}P\left[\phi\right]\left(t\right)\,.$
@@
$\left[\hat{a}\left(p\right),\hat{b}\left(q\right)\right]$ $=$ $\left(2\pi\right)^{d}\delta\left(p-q\right)\,,$
@@ IV.2 Normal ordering
$\phi\left(t,x\right)$$\equiv$ $e^{H_{{\rm FP}}t}\phi\left(x\right)e^{-H_{{\rm FP}}t}$
$\bar{\phi}\left(t,x\right)$$\equiv$ $e^{H_{{\rm FP}}t}\bar{\phi}\left(x\right)e^{-H_{{\rm FP}}t}\,.$
@@
$T\phi\left(t_{2},x_{2}\right)\phi\left(t_{1},x_{1}\right)$ $=$ $G_{\phi\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{1},x_{1}\right)\phi\left(t_{2},x_{2}\right):$
$T\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $G_{\bar{\phi}\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right):\,,$
@@
$T\phi\left(t_{2},x_{2}\right)\phi\left(t_{1},x_{1}\right)$ $=$ $G_{\phi\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right)^{2}:+\cdots$
$T\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $G_{\bar{\phi}\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right)\bar{\phi}\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right):+\cdots\,.$
It should be noted that, strictly speaking, the normal ordering represents one possible regularization for the composite operators of the free theory. In particular, the expectation value on any normal ordered operator vanishes by construction. However, this is just a possible choice and other choices, such as dim-reg discussed in section III.1, lead to different values of the VEV. Clearly, also the associated OPE coefficients are scheme dependent in general. However, one can obtain universal and scheme independent results, as we shall check explicitly at the fixed point.
@@
$T\bar{\phi}\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $:\bar{\phi}\left(\frac{t_{1}+t_{2}}{2},\frac{x_{1}+x_{2}}{2}\right)^{2}:+\cdots\,.$
@@ V Leading corrections to the OPE
One may adopt various strategies to compute OPE coefficients: for instance, one may look at the three- or four-point function and consider the short distance limit of two of the operators appearing in such correlations, see, e.g., Collins (2023). In this section we follow a different approach, based on the mass expansion of the two-point function, which leads to the same result, as we have checked explicitly. On the one hand, this approach allows one to derive the OPE coefficients and at the same time to calculate the two-point function close to criticality, i.e., for values of $m$ small but non-zero. Indeed, one of the first applications of the OPE was precisely the prediction of the form of the static two-point function close to criticality in the regime $m\ll p$ Brezin et al. (1974); Brézin et al. (1974); Caselle et al. (2015).
@@ V.1 OPE expansion of $\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$
In section III we have computed the VEV of the operator $\left[\phi^{2}/2\right]$ up to order $O\left(g\right)$. In order to apply the strategy outlined at the beginning of this section, we must compute the LHS of equation (15) up to the same order. At the tree level the two-point function is given by equation (23). The diagrams contributing to order $O\left(g\right)$ only involve the mass renormalization. Thus, it turns out that the LHS of (15) is given precisely by (23) once the substitution
@@
$c_{\phi\phi 1}^{\left(0\right)}\left(t,x\right)$ $=$ $\frac{2^{d-2}\Gamma\left(\frac{d}{2}-1\right)}{\left(4\pi\right)^{d/2}}\frac{1}{x^{d-2}}\Biggr{\{}1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\Biggr{\}}+O\left(\epsilon^{2}\right)$
$c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)$ $=$ $2+g_{*}\frac{1}{16\pi^{2}}\Biggr{\{}\log\left(\pi x^{2}\mu\right)+\gamma+\Gamma\left(0,\frac{x^{2}}{4Dt}\right)\Biggr{\}}+O\left(\epsilon^{2}\right)\,,$
@@ V.2 OPE expansion of $\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$
$\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$$\approx$ $\left(c_{\phi\bar{\phi}1}^{\left(0\right)}\left(t,x\right)+m^{2}c_{\phi\bar{\phi}1}^{\left(1\right)}\left(t,x\right)+\cdots\right)+\left(c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)+\cdots\right)\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}$
@@
$c_{\phi\bar{\phi}1}^{\left(0\right)}\left(t,x\right)$ $=$ $\frac{\theta\left(t\right)}{\left(4\pi\right)^{d/2}}\frac{e^{-\frac{x^{2}}{4Dt}}}{\left(Dt\right)^{d/2}}+O\left(\epsilon^{2}\right)$
$c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)$ $=$ $-g_{*}\frac{\theta\left(t\right)}{\left(4\pi\right)^{2}}\frac{e^{-\frac{x^{2}}{4Dt}}}{\left(Dt\right)}+O\left(\epsilon^{2}\right)\,,$
@@ V.3 Relation to the fluctuation-dissipation theorem
Again, we may consider the above relation off-criticality and consider the $m^{2}$-expansion. The leading term in this expansion is determined by the OPE coefficients associated with the identity, which must satisfy
$c_{\phi\bar{\phi}1}\left(t,x\right)$ $=$ $-\frac{1}{D}\partial_{t}c_{\phi\phi 1}\left(t,x\right)\,.$
@@
$c_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)$ $=$ $-\frac{1}{D}\partial_{t}c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)\,.$
It is straightforward to check that the coefficients calculated in sections V.1 and V.2 satisfy the above non-perturbative relations.
@@ VI Universal quantities from the OPE
In the previous section we obtained certain OPE coefficients. Now we shall see that they entail universal information after implementing suitable normalization conditions. The arguments used in this section follow the line of reasoning detailed in ref. Amit (1984), to which we refer for a review of universality in the context of the perturbative RG and its application to the statistical field theory approach for critical phenomena.
@@ VI.1 Universality in the static limit
$\frac{\langle\phi\left(p_{1}\right)\phi\left(p_{2}\right)\left[\frac{\phi^{2}}{2}\right]\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi\left(p_{1}-p_{2}\right)\phi\left(-p_{1}+p_{2}\right)\rangle\langle\left[\frac{\phi^{2}}{2}\right]\left(p_{1}+p_{2}\right)\left[\frac{\phi^{2}}{2}\right]\left(-p_{1}-p_{2}\right)\rangle^{1/2}}\,.$
@@
$\frac{\langle\phi_{1}\left(p_{1}\right)\phi_{1}\left(p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi_{1}\left(p_{1}-p_{2}\right)\phi_{1}\left(-p_{1}+p_{2}\right)\rangle\langle\phi_{2}\left(p_{1}+p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle^{1/2}}\,.$
@@
$\left(\mu\partial_{\mu}+\beta\partial_{g}\right)\frac{\langle\phi_{1}\left(p_{1}\right)\phi_{1}\left(p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi_{1}\left(p_{1}-p_{2}\right)\phi_{1}\left(-p_{1}+p_{2}\right)\rangle\langle\phi_{2}\left(p_{1}+p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle^{1/2}}$ $=$ $0\,.$
By using the arguments detailed in Amit (1984), one can see that the above ratio is universal. Thus, by applying the OPE to the three-point function above it is straightforward to conclude that the coefficient $c_{112}$ is universal too.
@@ VI.2.1 General discussion
$C\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta}}F\left(c_{1}\frac{t}{x^{z}}\right)\,,$
@@
$C\left(t,x\right)$ $=$ $\frac{c_{2}}{x^{\Delta}}f\left(c_{1}{\cal N}_{y}^{-1}y\right)\,\equiv\,\frac{c_{2}}{x^{\Delta}}f_{y}\left(y\right)\,,$
@@
$f_{y}\left(y=1\right)$ $=$ ${\rm const}.$
@@
$\langle\phi_{1}\left(0,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\,.$
@@
$C_{ijk}\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta_{i}+\Delta_{j}-\Delta_{k}}}c_{ijk}\left(y\right)\,,$
@@ VI.2.2 Application to our results
$\langle\phi_{1}\left(0,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\,,$
@@
${\cal N}_{1}$ $=$ $\left(\frac{2^{d-2}\Gamma\left(\frac{d}{2}-1\right)}{\left(4\pi\right)^{d/2}}\right)^{-1/2}\,.$
@@
$\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\Biggr{\{}1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\Biggr{\}}\,.$
@@
$x^{2\Delta_{\phi}}\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle\Bigr{|}_{y=1}$ $=$ $1-\frac{\Gamma\left(\frac{d}{2}-1,1\right)}{\Gamma\left(\frac{d}{2}-1\right)}\,.$
@@
$\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}g\left(y\right)\mbox{ with }g\left(y\right)=1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{1}{y}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\,,$
@@
$\langle\phi_{1}\left(t>0,x\right)\bar{\phi}_{1}\left(0,0\right)\rangle\Bigr{|}_{y=1}$ $=$ $\frac{1}{x^{d}}\,,$
@@
$\phi_{1}\times\phi_{1}$ $=$ ${\cal N}_{1}^{2}\phi\times\phi\,=\,{\cal N}_{1}^{2}c_{\phi\phi 1}+{\cal N}_{1}^{2}c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left[\frac{\phi^{2}}{2}\right]+\cdots$
$=$ ${\cal N}_{1}^{2}c_{\phi\phi 1}+{\cal N}_{1}^{2}c_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}{\cal N}_{2}^{-1}\phi_{2}+\cdots$
$=$ $c_{\phi_{1}\phi_{1}1}+c_{\phi_{1}\phi_{1}\phi_{2}}\phi_{2}+\cdots\,.$
@@
$c_{\phi_{1}\phi_{1}1}\left(t,x\right)$ $=$ $\frac{1}{x^{2\Delta_{\phi_{1}}}}\left\{1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{1}{y}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\right\}+O\left(\epsilon^{2}\right)$
$c_{\phi_{1}\phi_{1}\phi_{2}}\left(t,x\right)$ $=$ $\frac{1}{x^{2\Delta_{\phi_{1}}-\Delta_{\phi_{2}}}}\sqrt{2}\left(1-\frac{\epsilon}{6}+\frac{\epsilon}{6}\Gamma\left(0,\frac{1}{y}\right)\right)+O\left(\epsilon^{2}\right)\,.$
In the present approximation $c_{\phi_{1}\phi_{1}1}\left(t,x\right)$ is fixed by the Gaussian two-point function together with the chosen normalization conventions. The coefficient $c_{\phi_{1}\phi_{1}\phi_{2}}\left(t,x\right)$ is more interesting: in the static limit, i.e., $y\rightarrow 0$, we recover the $\epsilon$-correction known from CFT techniques Gaiotto et al. (2014); Gopakumar et al. (2017a, b); Dey et al. (2017); Gopakumar and Sinha (2018). In the dynamical case, this universal coefficient is promoted to be a function of the scaling variable $y$, which begins to be non-trivial from $O\left(\epsilon\right)$.
@@
$c_{\phi_{1}\bar{\phi}_{1}1}\left(t,x\right)$ $=$ $\frac{\theta(y)}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}}}\frac{4e^{-1/y}y^{-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}+O\left(\epsilon^{2}\right)$
$c_{\phi_{1}\bar{\phi}_{1}\phi_{2}}\left(t,x\right)$ $=$ $\frac{\theta(y)}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}-\Delta_{\phi_{2}}}}\left(-\epsilon\frac{e^{1-\frac{1}{y}}y^{-1}}{3\sqrt{2}}\right)+O\left(\epsilon^{2}\right)\,.$

### v2

@@ IV.1 Gaussian Fokker-Planck Hamiltonian
$\hat{\phi}\left(x\right)$ $=$ $\phi\left(x\right)$ (22)
$\hat{\pi}\left(x\right)$ $=$ $\frac{1}{i}\frac{\delta}{\delta\phi\left(x\right)}\,.$ (23)
@@
$\partial_{t}P\left[\phi\right]\left(t\right)$ $=$ $D\Biggr{\{}\int_{x}i\hat{\pi}\left(x\right)\left(-\partial_{x}^{2}+m^{2}\right)\hat{\phi}\left(x\right)-\int_{x}\hat{\pi}\left(x\right)^{2}\Biggr{\}}P\left[\phi\right]\left(t\right)\,.$ (24)
@@
$\left[\hat{a}\left(p\right),\hat{b}\left(q\right)\right]$ $=$ $\left(2\pi\right)^{d}\delta\left(p-q\right)\,,$ (27)
@@ IV.2 Normal ordering
$\phi\left(t,x\right)$$\equiv$ $e^{H_{{\rm FP}}t}\phi\left(x\right)e^{-H_{{\rm FP}}t}$ (29)
$\bar{\phi}\left(t,x\right)$$\equiv$ $e^{H_{{\rm FP}}t}\bar{\phi}\left(x\right)e^{-H_{{\rm FP}}t}\,.$ (30)
@@
$T\phi\left(t_{2},x_{2}\right)\phi\left(t_{1},x_{1}\right)$ $=$ $G_{\phi\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{1},x_{1}\right)\phi\left(t_{2},x_{2}\right):$ (31)
$T\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)$ $=$ $G_{\bar{\phi}\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right):\,,$ (32)
@@
$\boxed{T\phi\left(t_{2},x_{2}\right)\phi\left(t_{1},x_{1}\right)=G_{\phi\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{c},x_{c}\right)^{2}:+\cdots}$ (33)
and
$\boxed{T\phi\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)=G_{\bar{\phi}\phi}\left(t_{2}-t_{1},x_{2}-x_{1}\right)+:\phi\left(t_{c},x_{c}\right)\bar{\phi}\left(t_{c},x_{c}\right):+\cdots\,,}$ (34)
where $t_{c}\equiv\left(t_{1}+t_{2}\right)/2$ and $x_{c}\equiv\left(x_{1}+x_{2}\right)/2$. It should be noted that, strictly speaking, the normal ordering represents one possible regularization for the composite operators of the free theory. In particular, the expectation value on any normal ordered operator vanishes by construction. However, this is just a possible choice and other choices, such as dim-reg discussed in section III.1, lead to different values of the VEV. Clearly, also the associated OPE coefficients are scheme dependent in general. However, one can obtain universal and scheme independent results, as we shall check explicitly at the fixed point.
@@
$\boxed{T\bar{\phi}\left(t_{2},x_{2}\right)\bar{\phi}\left(t_{1},x_{1}\right)=:\bar{\phi}\left(t_{c},x_{c}\right)^{2}:+\cdots\,.}$ (35)
@@ V Leading corrections to the OPE
One may adopt various strategies to compute OPE coefficients: for instance, one may look at the three- or four-point function and consider the short distance limit of two of the operators appearing in such correlations, see, e.g., Collins (2023). In this section we follow a different approach, based on the mass expansion of the two-point function, which leads to the same result, as we have checked explicitly. On the one hand, this approach allows one to derive the OPE coefficients and at the same time to calculate the two-point function close to criticality, i.e., for values of $m$ small but non-zero. Indeed, one of the first applications of the OPE was precisely the prediction of the form of the static two-point function close to criticality in the regime $m\ll p$ Brezin et al. (1974); Brézin et al. (1974); Caselle et al. (2015). For a modern use of this idea in the static case see Caselle et al. (2015); Costagliola (2016); Caselle et al. (2016). Let us mentioned, however, that we checked that the results of this section can also be obtained by looking at the three- and four-point functions. In this way one has a non-trivial check of the consistency of the computed OPE coefficients.
@@ V.1 OPE expansion of $\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$
In section III we have computed the VEV of the operator $\left[\phi^{2}/2\right]$ up to order $O\left(g\right)$. In order to apply the strategy outlined at the beginning of this section, we must compute the LHS of equation (36) up to the same order. At the tree level the two-point function is given by equation (A). The diagrams contributing to order $O\left(g\right)$ only involve the mass renormalization. Thus, it turns out that the LHS of (36) is given precisely by (A) once the substitution
@@
$C_{\phi\phi 1}^{\left(0\right)}\left(t,x\right)$ $=$ $\frac{2^{d-2}\Gamma\left(\frac{d}{2}-1\right)}{\left(4\pi\right)^{d/2}}\frac{1}{x^{d-2}}\Biggr{\{}1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4D|t|}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\Biggr{\}}+O\left(\epsilon^{2}\right)$ (38)
$C_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)$ $=$ $2+g_{*}\frac{1}{16\pi^{2}}\Biggr{\{}\log\left(\pi x^{2}\mu\right)+\gamma+\Gamma\left(0,\frac{x^{2}}{4D|t|}\right)\Biggr{\}}+O\left(\epsilon^{2}\right)\,,$ (39)
@@ V.2 OPE expansion of $\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$
$\langle\phi\left(t,x\right)\bar{\phi}\left(0,0\right)\rangle$$\approx$ $\left(C_{\phi\bar{\phi}1}^{\left(0\right)}\left(t,x\right)+m^{2}C_{\phi\bar{\phi}1}^{\left(1\right)}\left(t,x\right)+\cdots\right)+\left(C_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)+\cdots\right)\Bigr{\langle}\left[\frac{\phi^{2}}{2}\right]\Bigr{\rangle}$ (40)
@@
$C_{\phi\bar{\phi}1}^{\left(0\right)}\left(t,x\right)$ $=$ $\frac{\theta\left(t\right)}{\left(4\pi\right)^{d/2}}\frac{e^{-\frac{x^{2}}{4Dt}}}{\left(Dt\right)^{d/2}}+O\left(\epsilon^{2}\right)$ (41)
$C_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}^{\left(0\right)}\left(t,x\right)$ $=$ $-g_{*}\frac{\theta\left(t\right)}{\left(4\pi\right)^{2}}\frac{e^{-\frac{x^{2}}{4Dt}}}{\left(Dt\right)}+O\left(\epsilon^{2}\right)\,,$ (42)
@@ V.3 Relation to the fluctuation-dissipation theorem
Again, we may consider the above relation off-criticality and consider the $m^{2}$-expansion. The leading term in this expansion is determined by the OPE coefficients associated with the identity, which, in the small mass limit, must satisfy
$\boxed{C_{\phi\bar{\phi}1}\left(t,x\right)=-\frac{1}{D}\partial_{t}C_{\phi\phi 1}\left(t,x\right)\,.}$ (45)
@@
$\boxed{C_{\phi\bar{\phi}\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)=-\frac{1}{D}\partial_{t}C_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left(t,x\right)\,}$ (46)
in the small mass limit. It is straightforward to check that the coefficients calculated in sections V.1 and V.2 satisfy the above non-perturbative relations.
@@ VI Universal quantities from the OPE
In the previous section we obtained certain OPE coefficients. Now we shall see that they entail universal information after implementing suitable normalization conditions. The arguments used in this section follow the line of reasoning detailed in Amit (1984), to which we refer for a review of universality in the context of the perturbative RG and its application to the statistical field theory approach for critical phenomena.
@@ VI.1 Universality in the static limit
$\frac{\langle\phi\left(p_{1}\right)\phi\left(p_{2}\right)\left[\frac{\phi^{2}}{2}\right]\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi\left(p_{1}-p_{2}\right)\phi\left(-p_{1}+p_{2}\right)\rangle\langle\left[\frac{\phi^{2}}{2}\right]\left(p_{1}+p_{2}\right)\left[\frac{\phi^{2}}{2}\right]\left(-p_{1}-p_{2}\right)\rangle^{1/2}}\,.$ (47)
@@
$\frac{\langle\phi_{1}\left(p_{1}\right)\phi_{1}\left(p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi_{1}\left(p_{1}-p_{2}\right)\phi_{1}\left(-p_{1}+p_{2}\right)\rangle\langle\phi_{2}\left(p_{1}+p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle^{1/2}}\,.$ (48)
@@
$\left(\mu\partial_{\mu}+\beta\partial_{g}\right)\frac{\langle\phi_{1}\left(p_{1}\right)\phi_{1}\left(p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle}{\langle\phi_{1}\left(p_{1}-p_{2}\right)\phi_{1}\left(-p_{1}+p_{2}\right)\rangle\langle\phi_{2}\left(p_{1}+p_{2}\right)\phi_{2}\left(-p_{1}-p_{2}\right)\rangle^{1/2}}$ $=$ $0\,.$ (49)
By using the arguments detailed in Amit (1984), one can see that the above ratio is universal. Thus, by applying the OPE to the three-point function above (i.e., taking a suitable large momentum limit) it is straightforward to conclude that the coefficient $c_{112}$ is universal too.
@@ VI.2.1 General discussion
$C\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta}}F\left(c_{1}\frac{t}{x^{z}}\right)\,,$ (50)
@@
$C\left(t,x\right)$ $=$ $\frac{c_{2}}{x^{\Delta}}f\left(c_{1}{\cal N}_{y}^{-1}y\right)\,\equiv\,\frac{c_{2}}{x^{\Delta}}f_{y}\left(y\right)\,,$ (52)
@@
$f_{y}\left(y=1\right)$ $=$ ${\rm const}.$ (53)
@@
$\langle\phi_{1}\left(0,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\,.$ (55)
@@
$C_{ijk}\left(t,x\right)$ $=$ $\frac{1}{x^{\Delta_{i}+\Delta_{j}-\Delta_{k}}}c_{ijk}\left(y\right)\,,$ (56)
@@ VI.2.2 Application to our results
$\langle\phi_{1}\left(0,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\,,$ (57)
@@
${\cal N}_{1}$ $=$ $\left(\frac{2^{d-2}\Gamma\left(\frac{d}{2}-1\right)}{\left(4\pi\right)^{d/2}}\right)^{-1/2}\,.$ (58)
@@
$\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}\Biggr{\{}1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4D|t|}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\Biggr{\}}\,.$ (59)
@@
$x^{2\Delta_{\phi}}\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle\Bigr{|}_{y=1}$ $=$ $1-\frac{\Gamma\left(\frac{d}{2}-1,1\right)}{\Gamma\left(\frac{d}{2}-1\right)}\,.$ (60)
@@
$\langle\phi_{1}\left(t,x\right)\phi_{1}\left(0,0\right)\rangle$ $=$ $\frac{1}{x^{2\Delta_{\phi}}}g\left(y\right)\mbox{ with }g\left(y\right)=1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{1}{|y|}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\,,$ (61)
@@
$\langle\phi_{1}\left(t>0,x\right)\bar{\phi}_{1}\left(0,0\right)\rangle\Bigr{|}_{y=1}$ $=$ $\frac{1}{x^{d}}\,,$ (62)
@@
$\phi_{1}\times\phi_{1}$ $=$ ${\cal N}_{1}^{2}\phi\times\phi\,=\,{\cal N}_{1}^{2}C_{\phi\phi 1}+{\cal N}_{1}^{2}C_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}\left[\frac{\phi^{2}}{2}\right]+\cdots$ (63)
$=$ ${\cal N}_{1}^{2}C_{\phi\phi 1}+{\cal N}_{1}^{2}C_{\phi\phi\left[\frac{\phi^{2}}{2}\right]}{\cal N}_{2}^{-1}\phi_{2}+\cdots$ (64)
$=$ $C_{\phi_{1}\phi_{1}1}+C_{\phi_{1}\phi_{1}\phi_{2}}\phi_{2}+\cdots\,.$ (65)
@@
$C_{\phi_{1}\phi_{1}1}\left(t,x\right)$ $=\frac{1}{x^{2\Delta_{\phi_{1}}}}\left\{1-\frac{\Gamma\left(\frac{d}{2}-1,\frac{1}{|y|}\right)}{\Gamma\left(\frac{d}{2}-1\right)}\right\}+O\left(\epsilon^{2}\right)$ (66a)
$C_{\phi_{1}\phi_{1}\phi_{2}}\left(t,x\right)$ $=\frac{1}{x^{2\Delta_{\phi_{1}}-\Delta_{\phi_{2}}}}\sqrt{2}\left(1-\frac{\epsilon}{6}+\frac{\epsilon}{6}\Gamma\left(0,\frac{1}{|y|}\right)\right)+O\left(\epsilon^{2}\right)\,.$ (66b)
In the present approximation $C_{\phi_{1}\phi_{1}1}\left(t,x\right)$ is fixed by the Gaussian two-point function together with the chosen normalization conventions. The coefficient $C_{\phi_{1}\phi_{1}\phi_{2}}\left(t,x\right)$ is more interesting: in the static limit, i.e., $y\rightarrow 0$, we recover the $\epsilon$-correction known from CFT techniques Gaiotto et al. (2014); Gopakumar et al. (2017a, b); Dey et al. (2017); Gopakumar and Sinha (2018). In the dynamical case, this universal coefficient is promoted to be a function of the scaling variable $y$, which begins to be non-trivial from $O\left(\epsilon\right)$. Let us note that we expect the equal-time limit of the OPE coefficient to coincide with the associated static coefficients. In this limit, non-perturbative values have been obtained by a variety of methods, such as conformal bootstrap Kos et al. (2016), Monte Carlo simulations Caselle et al. (2015); Costagliola (2016); Caselle et al. (2016), and functional renormalization group Rose et al. (2022). We refer to Rose et al. (2022) for a comparison of numerical results of the various approaches.
@@
$C_{\phi_{1}\bar{\phi}_{1}1}\left(t,x\right)$ $=\frac{\theta(y)}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}}}\frac{4e^{-1/y}y^{-d/2}}{\Gamma\left(\frac{d-2}{2}\right)}+O\left(\epsilon^{2}\right)$ (67a)
$C_{\phi_{1}\bar{\phi}_{1}\phi_{2}}\left(t,x\right)$ $=\frac{\theta(y)}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}-\Delta_{\phi_{2}}}}\left(-\epsilon\frac{e^{1-\frac{1}{y}}y^{-1}}{3\sqrt{2}}\right)+O\left(\epsilon^{2}\right)\,.$ (67b)